# Exploring Influence Factors on LLM Suitability for No-Code Development of End User IoT Applications

### v1

@@ 3.2.1. Model Selection
With the rapid proliferation of LLMs, it is essential to evaluate their feasibility within no-code development platforms. A thorough analysis of various LLMs can help optimize model selection strategies, ensuring these platforms achieve peak performance, efficiency, and user-friendliness. To explore this, we compare five mainstream LLMs from different aspects of consideration, i.e.,
@@
This comparison provides insights into the trade-offs between model size, design focus, and domain-specific optimization, guiding future adoption strategies for LLM-powered no-code platforms.
@@ 3.2.2. Prompt Language
The input language of LLMs can presumably have a large impact depending on the text corpora a model is trained on. Models developed in English-speaking countries, i.e., English models, will presumably be trained with more English texts while models developed in from Chinese developer teams, Chinese models, may handle Chinese user input better. To assess this language-specific impact, we conduct an experiment examining how prompt language, i.e., English and Chinese, affects LLMs performance.
@@ 3.2.3. Community Background
The consideration of prompt language choice also underscores the importance to evaluate the performance with Chinese LLMs. We select three mainstream Chinese LLMs to evaluate their performance, i.e., Alibaba Qwen, Baidu Qianfan and DeepSeek R1, which also differ in design focus.
@@
    Alibaba Qwen, a model leveraging a Mixture-of-Experts (MoE) architecture, which good at handling long-context tasks and large-scale language understanding.
@@
By comparing the performance of Chinese and English LLMs, we investigate whether aligning an LLMs training corpus with the input language improves its performance of LLM-powered no-code platforms.
@@ 3.3. Dataset
Figure 1. Dataset Example: We use a LLM-powered No-Code Platform, i.e., LLM4FaaS, as the base no-code platform for evaluation. Specifically, it takes the user natural language description in combined with the project context and reference code as input to LLMs. To evaluate the accuracy of LLM-generated results, we set the ground truth based on the user requirements. The accuracy is set by comparing the output from generated results and the ground truth. Essential, the dataset consists of user requirements and the corresponding ground truth.
We use and extent an existing dataset of 26 real users natural language description for 4 smart home automation tasks in varying levels of complexity(Wang et al., 2025). Specifically, the complexity of task from simple to complex varies as follows and present an example in Figure 2:
@@
    Simple: a use case which expecting a straightforward device control function.
@@
    Medium: introduce an additional layer of complexity, where the system is expected to handle three keyword-based or time-based automation tasks.
@@
    Advanced: the scenario complexity improves by triggering smart devices based on the real-time sensor readings.
@@
Figure 2. User Response Example: This figure presents a real user response from the dataset, illustrating four smart home automation tasks across different complexity levels. The response showcases increasing complexity, from simple device control to intricate automation involving real-time sensor data and user-environment coordination. The original response is in Chinese, and we provide an English translation for clarity.
@@
Input Language Variation
To assess the impact of input language on LLMs performance, we translate the original Chinese user responses into English using Google Translate [^2] https://translate.google.com/, as English is the primary language for most LLMs. This allows for a direct comparison of LLM outputs across different input languages, i.e., Chinese and English, revealing potential language biases and inconsistencies.
Ground Truth Definition
To evaluate the performance of LLM-generated results, we define the ground truth based on user responses. Specifically, we construct the ground truth based on the expected console output, derived from user responses and the intended system response. The ground truth is determined on a per-user and per-task basis, representing the expected system response for each automation task from every user. This ensures consistency in evaluation and allows for an objective comparison between LLM-generated outputs and the expected results.
@@ 3.4. Evaluation Metrics
To address this limitation, we refine the evaluation metrics that quantify semantic correctness on a continuous scale, allowing for a more granular assessment of LLM-generated outputs. In detail, the semantic accuracy rate use coverage match rate (CMR) to measure the extent to which the ground truth (GT) outputs are successfully covered by the LLM-generated outputs (MO), allowing for variations in wording and additional non-disruptive information in MO. CMR quantifies the proportion of GT entries that are correctly identified within MO. It is formally defined as:
@@ 4. Results
We present the experimental results, analyzing the impact of key factors on the performance of LLM-powered no-code platforms. The results highlight key findings across model selection (Section 4.1), prompt language (Section 4.2), and community background (Section 4.3), providing insights into their respective impacts.
@@ 4.1. Model Selection
The ranking of model performances for all or most experiments remains stable in the following order: GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Copilot, and Llama-3.1. Among these, GPT-4o and GPT-4o-mini show the best performance in both syntactic and semantic success. Gemini and Copilot have a similar performance for average semantic and syntactic success rate, Copilot performs better on simple tasks, and Gemini in complex tasks. Llama fails to generate functions with the user prompts. Specifically, the average syntactic success rate of GPT-4o and GPT-4o-mini is 89.10% and 73.99%, respectively, while the other LLMs are below 50%. The average semantic success rate of GPT-4o is 67.54%, whereas the second best, GPT-4o-mini, has dropped to 32.07%, and other LLMs is below 15.00%. In addition to the semantic success rate, we are also curious about how close the results are to the semantic success, given by the semantic accuracy rate. Figure 4 depicts the semantic accuracy, showing a high-level of semantic accuracy of GPT-4o which still demonstrates a clear advantage over the competing models across all difficulty levels. For all the models, the distribution of semantic accuracy of results becomes more dispersed as the task complexity increases. Notably, the results of complex tasks sometimes exhibit higher semantic success than those of medium and advanced tasks, i.e., it does not always align with the predefined task complexity. This may be due to the fact that both medium and advanced tasks involve three sub-tasks that are unrelated to each other, which can lead to more semantic failures.
@@ 4.2. Prompt Language
Figure 5. Success rates with English prompts. GPT-4o still outperforms other models with English prompts in both syntactic and semantic accuracy.
@@
To investigate the impact of prompt language to LLM performance, we use the translated English user prompt from dataset and evaluate the performance among the five LLMs. We show the syntactic and semantic success rate of English prompt results in Figure 5, then show the comparison to original results in Figure 6 and Figure 7.
GPT-4o and GPT-4o-mini, i.e., OpenAI models maintain relatively high syntactic success rates, i.e. 89.78% and 71.12%, respectively, while the other LLMs remain below 50%. Llama still fails to generate functions for all the prompts, resulting in 0% syntactic and semantic success rates.
For GPT-4o, the syntactic success rate increases slightly by 0.69 percentage points, while the semantic success rate decreases by 12.34 percentage points compared to the original prompts. English prompts improve performance only on easy tasks but negatively impact more complex ones. The average syntactic and semantic success rates of GPT-4o-mini almost remain unchanged, with a 1.68 percentage point increase in semantic success rate and a 2.87 percentage point decrease in syntactic success rate. For tasks in different complexity levels, it shows an opposite trend to GPT-4o, where it drops for easy and medium tasks but improves for advanced and complex ones. The translation step helps GPT-4o-mini to better understand the user requirements with advanced and complex tasks, while losing nuances of the original user requirements leads to a performance drop in the easy and medium tasks. Gemini experiences a decrease in both syntactic and semantic success rates by 8.77 percentage points and 6.73 percentage points, respectively. Copilot, in contrast, demonstrates an overall improvement of 9.17 percentage points in syntactic success and 5.7 percentage points in semantic success. Specifically, with English prompts, Copilot achieves a comparable performance to GPT-4o-mini with the easy task.
@@ 4.3. Community Background
Figure 9. Semantic Accuracy Comparison with Mandarin Models: GPT-4o also outperforms the Mandarin models in semantic accuracy across tasks of varying difficulty, but the differences are less pronounced. The two Mandarin models showcase a similar performance of semantic accuracy where Baidu-Qianfan is slightly better. With the task complexity increase, the semantic accuracy distribution of results becomes more dispersed for all models.
@@
Considering LLMs trained with language-specific data may have a better understanding of the user prompts in the same language, we evaluate the performance of three Chinese LLMs, i.e., Baidu-Qianfan, Ali-Qwen-Max and DeepSeek-r1:7b. We compare their performance with the OpenAI models, which performed best in the initial experiments and show the results in Figure 8 and Figure 9.
@@
Baidu-Qianfan shows a same syntactic performance to GPT-4o with an 89.10% syntactic success rate, followed by Ali-Qwen-Max, GPT-4o-mini and DeepSeek with 77.79%, 73.99% and 46.92%, respectively. In terms of the overall semantic success rate, GPT-4o maintains its lead with 67.54%, outperforming the other models. Baidu-Qianfan, Ali-Qwen-Max, and GPT-4o-mini follow with success rates of 51.93%, 43.28%, and 32.07%, respectively. DeepSeek entirely fails to generate function that meets user request, lead to a 0% semantic success rate. Notably, Baidu-Qianfan can achieve a similar semantic success rate (78.26%) to GPT-4o (82.61%) in the easy task but not for more complex ones. Ali-Qwen-Max has a comparable performance to GPT-4o-mini in the easy task but shows a clear advantage in the tasks with higher complexity. Excluding DeepSeek which has a 0% semantic success rate, GPT-4o-mini has the worst performance among this comparison, which can be caused by both the model capability and understanding of prompt language.

### v2

@@ 3.2.1. Model Selection
With the rapid proliferation of LLMs, it is essential to evaluate their feasibility within NCDPs. A thorough analysis of various LLMs can help optimize model selection strategies, ensuring these platforms achieve peak performance, efficiency, and user-friendliness. We intentionally select models that vary in architecture, scale, and domain specialization to investigate how fundamental design properties affect model behavior. In this way, we can derive insights that are more robust and generalizable, beyond the short-term performance of any single model version.
To explore this, we compare five mainstream LLMs from different aspects of consideration, i.e.,
@@
This comparison provides insights into the trade-offs between model size, design focus, and domain-specific optimization, guiding future adoption strategies for LLM-powered NCDPs.
@@ 3.2.2. Prompt Language
The input language of LLMs can presumably have a large impact depending on the text corpora a model is trained on. Models developed in English-speaking countries, i.e., English models, will presumably be trained with more English texts while models developed in from Chinese developer teams, i.e., Chinese models, may handle Chinese user input better. To assess this language-specific impact, we conduct an experiment examining how prompt language, i.e., English and Chinese, affects LLMs performance.
@@ 3.2.3. Community Background
The consideration of prompt language choice also underscores the importance of evaluating the performance with Chinese LLMs. We select three mainstream Chinese LLMs to evaluate their performance, i.e., Alibaba Qwen, Baidu Qianfan and DeepSeek R1, which also differ in design focus.
@@
    Alibaba Qwen, a model leveraging a Mixture-of-Experts (MoE) architecture, which is good at handling long-context tasks and large-scale language understanding.
@@
By comparing the performance of Chinese and English LLMs, we investigate whether aligning an LLMs training corpus with the input language improves its performance of LLM-powered NCDPs.
3.2.4. Few-shot vs. Zero-shot Experiment Setting
To assess the impact of prompting strategies on the performance of LLM-based NCDPs, we additionally conduct a few-shot experiment and compare it with the default zero-shot setting. Specifically, we design a feedback-based few-shot experiment leveraging runtime syntactic error as dynamic, task-specific guidance for iterative code refinement. This approach avoids the potential noise and bias introduced by irrelevant examples, which are difficult to predefine given the highly customized and user-specific nature of NCDP tasks, especially for non-technical users.
  • Zero-shot setting: The LLMs generate code based solely on a structured prompt that contains the user requirements, without access to prior outputs or error feedback.
  • Few-shot setting: The LLM receives additional context in the form of runtime error messages from previous attempts. These serve as feedback to guide subsequent code refinement. The generation process is allowed to iterate up to three times.
The few-shot setup aims to simulate a realistic development scenario by iteratively refining code with the feedback of interpreter or compiler, providing insights into the self-correct ability of LLM. We restrict the experiment to syntactic errors, which yield objective and consistent outcomes. In contrast, semantic error handling typically requires clarification of user intent, which is inherently subjective, particularly in NCDP contexts involving non-technical users, and thus infeasible for standardized evaluation.
@@ 3.3. Dataset
We use and extend an existing dataset of 26 real users natural language description for 4 smart home automation tasks in varying levels of complexity(Wang et al., [n. d.]). Specifically, the complexity of task from simple to complex varies as follows, and we present an example in Figure 1.
@@
    Simple: a use case which expects a straightforward device control functionality.
@@
    Medium: introduce an additional layer of complexity, where the system is expected to handle three keyword-based or time-based automation sub-tasks.
@@
    Advanced: the scenario complexity improves by involving three sub-tasks and triggering smart devices based on the real-time sensor readings.
@@
Figure 1. User Response to Tasks of Varying Complexity: This figure presents a real user response from the dataset, illustrating four smart home automation tasks across different complexity levels. The response showcases increasing complexity, from simple device control to intricate automation involving real-time sensor data and user-environment coordination. The original response is in Chinese, and we provide an English translation for clarity.
Figure 2. Dataset Example: We use a LLM-powered No-Code Platform, i.e., LLM4FaaS, as the base NCDP for evaluation. Specifically, it takes the user natural language description in combined with the project context and reference code as input to LLMs. To evaluate the accuracy of LLM-generated results, we set the ground truth based on the user requirements. The accuracy is set by comparing the output from generated results and the ground truth. Essentially, the dataset consists of user requirements and the corresponding ground truth.
3.3.1. Input Language Variation
To assess the impact of input language on LLMs performance, we translate the original Chinese user responses into English using Google Translate (Google Translate, [n. d.]), as English is the primary language for most LLMs. This allows for a direct comparison of LLM outputs across different input languages, i.e., Chinese and English, revealing potential language biases and inconsistencies.
3.3.2. Ground Truth Definition
To evaluate the performance of LLM-generated results, we define the ground truth, constructed based on the expected console output, which is derived from user responses and the intended system response. The ground truth is determined on a per-user and per-task basis, representing the expected system response for each automation task from every user. This ensures consistency in evaluation and allows for an objective comparison between LLM-generated outputs and the expected results.
@@ 3.4. Evaluation Metrics
To address this limitation, we refine the evaluation metrics that quantify semantic accuracy on a continuous scale, allowing for a more granular assessment of LLM-generated outputs. In detail, the semantic accuracy rate use coverage match rate (CMR) to measure the extent to which the ground truth (GT) outputs are successfully covered by the LLM-generated outputs (MO), allowing for variations in wording and additional non-disruptive information in MO. CMR quantifies the proportion of GT entries that are correctly identified within MO. It is formally defined as:
@@
Our three metrics, i.e., syntactic success, semantic success, and semantic accuracy, are conceptually aligned with standard evaluation metrics such as precision, recall, and F1-score, but adapted for NCDPs. Specifically, syntactic success checks whether generated code runs without errors, i.e., a prerequisite for meaningful evaluation. Semantic success is analogous to precision: the proportion of syntactically valid outputs that fully satisfy the task requirements. Semantic accuracy is analogous to recall: the degree to which the task requirements are satisfied, even partially, by valid outputs. Taken together, semantic success and semantic accuracy jointly reflect both correctness and completeness, thus serving a role similar to the F1-score in capturing overall task-level performance. We adopt these tailored metrics to better reflect utility and reliability in no-code settings.
@@ 4. Results
We present the experimental results, analyzing the impact of key factors on the performance of LLM-powered NCDPs. The results highlight key findings across model selection (Section 4.1), prompt language (Section 4.2), community background (Section 4.3), and zero-shot vs. few-shot settings (Section 4.4), providing insights into their respective impacts (Section 4.5).
@@ 4.1. Model Selection
The ranking of model performances for all or most experiments remains stable in the following order: GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Copilot, and LLaMA-3.1. Among these, GPT-4o and GPT-4o-mini show the best performance in both average syntactic and semantic success rate among all tasks. Gemini and Copilot have a similar performance for average semantic and syntactic success rate, where Copilot performs better on simple tasks, and Gemini in complex tasks. LLaMA fails to generate functions with all the user prompts. Specifically, the average syntactic success rate among all task complexity levels of GPT-4o and GPT-4o-mini is 89.10% and 73.99%, respectively, while the other LLMs are below 50%. The average semantic success rate of GPT-4o among all tasks is 67.54%, whereas the second best, GPT-4o-mini, has dropped to 32.07%, and other LLMs is below 15.00%.
In addition to the semantic success rate, we are also curious about how close the results are to the semantic success, given by the semantic accuracy rate. Figure 4 depicts the semantic accuracy, showing a high-level of semantic accuracy of GPT-4o, which still demonstrates a clear advantage over the competing models across tasks in all difficulty levels. For all the models, the distribution of semantic accuracy rate becomes more dispersed as the task complexity increases. Notably, the results of complex tasks sometimes exhibit higher semantic success than those of medium and advanced tasks, i.e., it does not always align with the predefined task complexity. This may be due to the fact that both medium and advanced tasks involve three unrelated sub-tasks, which can lead to more semantic failures.
@@ 4.2. Prompt Language
Figure 5. Success rates with English prompts. GPT-4o still outperforms other models with English prompts in both syntactic and semantic success.
@@
To investigate the impact of prompt language to LLM performance, we use the translated English prompt from dataset and evaluate the performance among the five LLMs. We show the syntactic and semantic success rate of English prompt results in Figure 5, then show the comparison to original results in Figure 6 and Figure 7.
GPT-4o and GPT-4o-mini, i.e., OpenAI models, maintain relatively high average syntactic success rates, i.e. 89.78% and 71.12%, respectively, while the other LLMs remain below 50%. LLaMA still fails to generate functions for all the prompts, resulting in 0% syntactic and semantic success rates. For GPT-4o, the syntactic success rate increases slightly by 0.69 percentage points, while the semantic success rate decreases by 12.34 percentage points compared to the original prompts. English prompts improve performance only on easy tasks but negatively impact more complex ones. The average syntactic and semantic success rates of GPT-4o-mini almost remain unchanged, with a 1.68 percentage point increase in semantic success rate and a 2.87 percentage point decrease in syntactic success rate. For tasks in different complexity levels, it shows an opposite trend to GPT-4o, where it drops for easy and medium tasks but improves for advanced and complex ones. The translation step helps GPT-4o-mini to better understand the user requirements with advanced and complex tasks, while losing nuances of the original user requirements leads to a performance drop in the easy and medium tasks. Gemini experiences a decrease in both syntactic and semantic success rates by 8.77 percentage points and 6.73 percentage points, respectively. Copilot, in contrast, demonstrates an overall improvement of 9.17 percentage points in syntactic success and 5.7 percentage points in semantic success. Specifically, with English prompts, Copilot achieves a comparable performance to GPT-4o-mini with the easy task.
@@ 4.3. Community Background
Figure 9. Semantic Accuracy Comparison with Chinese Models: GPT-4o also outperforms the Chinese models in semantic accuracy across tasks of varying difficulty, but the differences are less pronounced. The two Chinese models showcase a similar performance of semantic accuracy where Baidu-Qianfan is slightly better. With the task complexity increase, the semantic accuracy distribution of results becomes more dispersed for all models.
@@
Considering LLMs trained with language-specific data may have a better understanding of the user prompts in the same language, we evaluate the performance of three Chinese LLMs, i.e., Baidu-Qianfan, Ali-Qwen-Max and DeepSeek-r1:7b, with original Chinese user requirements. We compare their performance with the OpenAI models, which performed best in the initial experiments and show the results in Figure 8 and Figure 9.
@@
Baidu-Qianfan shows a same syntactic performance to GPT-4o with an 89.10% syntactic success rate, followed by Ali-Qwen-Max, GPT-4o-mini and DeepSeek with 77.79%, 73.99% and 46.92%, respectively. In terms of the average semantic success rate of all tasks, GPT-4o maintains its lead with 67.54%. Baidu-Qianfan, Ali-Qwen-Max, and GPT-4o-mini follow with success rates of 51.93%, 43.28%, and 32.07%, respectively. DeepSeek entirely fails to generate function that meets user request, lead to a 0% semantic success rate. Notably, Baidu-Qianfan can achieve a similar semantic success rate (78.26%) to GPT-4o (82.61%) in the easy task but not for more complex ones. Ali-Qwen-Max has a comparable performance to GPT-4o-mini in the easy task, and shows a clear advantage in the tasks with higher complexity. Excluding DeepSeek which has a 0% semantic success rate, GPT-4o-mini has the worst performance among this comparison, which can be caused by both the model capability and understanding of prompt language.
4.4. Zero-shot vs. Few-shot Comparison
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 10. Syntactic and Semantic Success Rate after Few-Shot Experiment with all models. We show few-shot and zero-shot results using high and low opacity bars, respectively. For all models, giving error messages as feedback can improve both the syntactic and semantic success rate.
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 11. Semantic Accuracy of Few-Shot Experiment with Prompt in Chinese among all models. The semantic accuracy shows a similar bimodal distribution as the zero-shot setting results, while GPT-4o and Ali-Qwen-Max tends to be more concentrated around 100% semantic accuracy among all tasks.
To better understand the effect of iterative feedback on code generation performance, we compare the zero-shot and few-shot prompting strategies across multiple LLMs. This evaluation aims to quantify how compiler-level feedback, specifically, syntactic error messages, can help LLMs refine their outputs. In the few-shot experiments, we focus exclusively on syntactic errors, as these can be reliably and objectively detected by compilers. In contrast, semantic errors are inherently more difficult to define and evaluate, particularly in the NCDPs context, where non-technical users may express requirements in diverse and subjective ways. We evaluate the same set of LLMs as in previous experiments, including GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Copilot, Ali-Qwen-Max, and DeepSeek-r1:7b. We exclude Baidu-Qianfan and LLaMA from the few-shot experiment, as incorporating error messages as feedback to LLM would exceed the input token limit of Baidu-Qianfan, and LLaMA exhibits an all-zero performance in the zero-shot setting. Also, we use the English prompts in the few-shot experiment to showcase the input language influence.
4.4.1. Few-Shot: Model Selection
Providing error messages as feedback consistently enhances syntactic success across all models, with some achieving 100% syntactic success. This syntactic correction also leads to improvements in semantic success. We show the syntactic and semantic success rate of few-shot results in Figure 10, and present the semantic accuracy distribution in Figure 11.
The overall performance ranking in the few-shot setting differs slightly from the zero-shot setting. While GPT-4o remains the top model, GPT-4o-mini, which ranked second in the zero-shot experiments, drops to third place, overtaken by Copilot. In detail, GPT-4o and GPT-4o-mini, i.e., the top-performing models in the zero-shot setting, continue to show strong performance in the few-shot settings. Specifically, GPT-4o-mini achieves a 100% syntactic success rate for all tasks, while GPT-4o fails to resolve one syntactic error, resulting in a 95.65% syntactic success rate, all other tasks yield 100% syntactic success. Regarding semantic success rate, GPT-4o achieves an average of 76.57% across all tasks, while GPT-4o-mini has a 39.41% semantic success rate, comparing to the 67.54% and 32.07% in the zero-shot setting. Additionally, we measure the semantic accuracy, representing how close the model outputs are to semantic success. GPT-4o achieves an average of 87.65% across all tasks, while GPT-4o-mini has a 61.42% semantic accuracy, comparing to the 79.20% and 45.97% in the zero-shot setting.
In the few-shot setting, Copilot shows substantial improvements in both syntactic and semantic success rate, achieving an average syntactic success rate of 95.75% and an average semantic success rate of 47.25% across all tasks, comparing to 34.43% and 13.35%, respectively, in the zero-shot setting. Copilot does not achieve a 100% syntactic success rate as GPT-4o-mini, however, it surpasses it in terms of average semantic success rate. In particular, Copilot outperforms GPT-4o-mini on all but the easy task when measuring semantic success. Copilot tends to first generate a skeleton or pseudocode of the user required function in the first round, and fill the function logic in the following iterations, i.e., it acts more like a human developer. It demonstrates a 66.49% semantic accuracy rate in the few-shot setting, which is an improvement over the zero-shot setting at 19.78%.
Gemini showcases a syntactic improvement compared to the zero-shot setting, however, this enhancement has limited impact on its semantic performance. Except for the easy task, where Gemini achieves an average 39.13% semantic success rate, other tasks are around 10%.
4.4.2. Few-Shot: Community Background
We present the Chinese LLMs results alongside those of English LLMs results in Figure 10 and Figure 11 because the performance gap among the LLMs is narrowing. Considering all the LLMs evaluated in the few-shot setting, the performance ranking is GPT-4o, Ali-Qwen-Max, Copilot, GPT-4o-mini, Gemini, and DeepSeek R1.
In the few-shot setting, Ali-Qwen-Max continues to demonstrate strong performance and significant improvements. It achieves a syntactic success rate of 100% across all tasks, and an average semantic success rate of 57.93%, up from 77.79 and 43.28%, respectively, in the zero-shot setting. The semantic accuracy of Ali-Qwen-Max also improves to an average of 81.02%, compared to 63.54% in the zero-shot setting, narrowing the gap with GPT-4o at 87.65%. In the complex task, Ali-Qwen-Max outperforms GPT-4o in the semantic success rate, achieving 71.43% compared to GPT-4o’s 66.67%. While it can be partially attributed to a higher number of syntactic error cases, Ali-Qwen-Max still shows a strong performance in resolving syntactic errors and improving semantic success performance.
DeepSeek R1 also shows a syntactic improvement, increasing from 46.92% to 63.81%. However, it still fails to achieve any semantic success or semantic accuracy among all tasks.
4.4.3. Few-Shot: Prompt Language
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 12. Syntactic and Semantic Success Rate after Few-Shot Experiment with English prompts. Similar with using Chinese prompts, the few-shot experiments with English prompts also improves in syntactic and semantic success rate for all models.
Since prompt language can affect LLM performance, particularly when it aligns with the training data of LLMs, we also assess LLMs using English prompts under the few-shot setting. We show the syntactic and semantic success rate of few-shot results in Figure 12, and present the comparison with prompt in Chinese in Figure 13 and Figure 14.
The results indicate that all the models exhibit improvements in the syntactic success rate, but leading to different trend of improvement in the semantic success rate. For the OpenAI models, i.e., GPT-4o and GPT-4o-mini, the few-shot experiment with English prompts improves the syntactic success rate but slightly improve the semantic success rate. Specifically, for syntactic success rate, GPT-4o improves from an average of 89.78% to 100%, and GPT-4o-mini improves from 71.81% to 97.73%. In terms of semantic success rate, GPT-4o shows improvement only on the medium task, the others keep unchanged, leading to an average semantic success rate improvement of 2.28%. GPT-4o-mini shows an average semantic improvement of 5.36%, increasing from 33.97% to 39.33%, with no change only on the complex task.
Copilot, on the other hand, demonstrates a significant improvement in both syntactic and semantic success rate among all tasks, with an average syntactic success rate of 98.91% and an average semantic success rate of 52.01%, comparing to those in the zero-shot setting of 43.60% and 19.05%, respectively. Notably, Copilot outperforms GPT-4o for the advanced and complex tasks in the semantic success. This is partly due to the low number of syntactic error cases of GPT-4o in the few-shot setting, which limits opportunities for feedback-based improvement. It demonstrates the coding ability of Copilot with the pseudocode or code skeleton generated in the zero-shot experiment. For the semantic accuracy, Copilot achieves an average of 73.75% across all tasks, a significant improvement compared over its zero-shot results at 25.45%. While Gemini shows notable improvement in syntactic success rate, from 34.58% to 69.33%, its improved average semantic success rate remains low at 13.10%. Overall, using English prompts can help GPT-4o-mini and Copilot to achieve performance improvement, particularly in semantic understanding and code generation quality under the few-shot setting.
(a) GPT-4o
(b) GPT-4o-mini
(c) Gemini
(d) Copilot
Figure 13. Syntactic success rate comparison with prompts in Chinese and English under the few-shot setting. We denote Chinese and English prompts as ZH and EN in the figure, respectively. The language choice of the prompt has little impact on the syntactic success rate.
Comparing the results with those using Chinese prompts, we find that for GPT-4o, GPT-4o-mini and Copilot, iteratively incorporating error messages enables nearly 100% syntactic success rates. Although Gemini gains improvement, it still lags behind the others. For GPT-4o and GPT-4o-mini, the prompt language have little impact on the syntactic success. In contrast, both Copilot and Gemini syntactically perform better when using English prompts than Chinese prompts. As for the semantic success rate, using the Chinese prompts which containing the original user description performs better with GPT-4o and Gemini. Using English prompts helps Copilot achieve a better semantic success due to the elimination of translation step during the generation process. GPT-4o-mini has a better semantic success rate with Chinese prompts for the easy task, as the task complexity increases, the English prompts may help to better understand the user intents.
(a) GPT-4o
(b) GPT-4o-mini
(c) Gemini
(d) Copilot
Figure 14. Semantic success rate comparison with Chinese and English prompts under the few-shot setting. We denote Chinese and English prompts as ZH and EN, respectively. GPT-4o generally shows outstanding performance in semantic success rate under both Chinese and English prompts with few-shot setting.
4.4.4. Iteration Required to Fix Syntactic Errors
(a) Experiment with Chinese Prompts
(b) Experiment with English Prompts
Figure 15. We show the amount of syntactic error cases of zero-shot experiment results as Initial Errors for the few-shot setting. The Fixed Errors bar shows the number of cases leads to a syntactic success within a maximum of 3 iterations.
(a) Experiment with Chinese Prompts
(b) Experiment with English Prompts
Figure 16. We show the iterations needed to fix syntactic errors for the experiments with Chinese and English prompts among tasks in different levels of complexity. Zero represents initial syntactic error cases from zero-shot results.
In the few-shot experiments, we allow each LLM to iterate up to three times to resolve syntactic errors observed in the zero-shot results, and record the number of syntactic error cases that can be successfully corrected within these rounds. We show the results in Figure 15 and Figure 16.
When using Chinese prompts, both GPT-4o-mini and Ali-Qwen-Max successfully resolved all syntactic errors within three iterations. GPT-4o and Copilot corrected 90.00% and 93.44% of syntactic errors, respectively. Specifically, GPT-4o failed to resolve 1 out of 10 error cases, while Copilot failed in 4 out of 61 cases. At the lower end of performance, Gemini and DeepSeek R1 corrected only 44.44% and 32.65% of syntactic errors, respectively. When using English prompts, GPT-4o resolves all syntactic errors, followed by Copilot, GPT-4o-mini and Gemini with 98.11%, 96.15%, and 53.33% syntactic success rate, respectively.
While achieving high syntactic error fixing rates, the number of iterations required varies. Using Chinese prompts, GPT-4o resolves all syntactic errors in the first iteration. GPT-4o-mini and Ali-Qwen-Max can fix all syntactic errors within two rounds, i.e., around 90% in the first iteration, and around 10% in the second iteration. With English prompts, GPT-4o and GPT-4o-mini fix errors within two iterations. Specifically, GPT-4o required a second round for 1 out of 9 cases. Copilot demonstrates strong performance by fixing nearly all syntactic errors in the first iteration, with only one case carried over to the second and one additional case requiring a third round out of 52 cases.